# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ 7 Conclusion
Linear Gaussian inverse problems on possibly infinite-dimensional Hilbert spaces are an important kind of nonparametric inverse problem. For example, they can be used to approximate nonlinear nonparametric problems using the Laplace approximation, and they often serve as the native infinite-dimensional formulation of linear inverse problems before the parameter space $\mathcal{H}$ is discretised and they are in this sense ‘dimension independent’.
Optimal low-rank approximation of the posterior covariance for a class of losses that includes the KL divergence and the Hellinger metric, and of the posterior mean for the Bayes Risk, was studied in [30]. The analysis showed that certain matrix pencils, namely the ones defined by the Hessian and prior covariance and the prior and posterior covariance, form the central objects of study.
So far, these results applied to finite-dimensional parameter spaces only. In this work we have derived optimal low-rank approximations to the posterior covariance, posterior mean, and entire posterior distribution, on possibly infinite-dimensional parameter spaces, thereby generalising the conclusions of [30]. The resulting posterior approximations are equivalent with the exact posterior distribution. Moreover, the optimal low-rank updates for the covariance and mean share the property with the exact updates that their range lies in $\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$. We have shown that the posterior covariance approximations are optimal for a class of losses which includes the KL divergence, the reverse KL divergence, the Hellinger metric and the Rényi divergences. The constructed low-rank approximation of the mean is optimal for the reverse KL divergence and the Rényi divergences, and the low-rank approximation of the mean and covariance jointly is optimal for the reverse KL divergence. Furthermore, we have shown that the operator pencils which proved central in the finite-dimensional analysis, are both equivalent representations of the Hilbert–Schmidt operator appearing in the Feldman–Hajek theorem which quantifies similarity of Gaussian measures. It is therefore this operator that is central to the reduction of the posterior distribution for linear Gaussian inverse problems.
Finally, the low-rank approximations constructed in this work, which have been proven to be optimal in various measures of divergence, may provide a basis for the development of computationally efficient approximation of certain linear Gaussian problems, such as Gaussian Process Regression, and for optimal approximation of nonlinear nonparametric inverse problems.
@@ 8 Acknowledgements
The research of the authors has been partially funded by the Deutsche Forschungsgemeinschaft (DFG) Project-ID 318763901 – SFB1294. The authors thank Ricardo Baptista and Youssef Marzouk for helpful remarks concerning the joint approximation problem, and Bernhard Stankewitz for helpful discussions.
@@ Appendix A Auxiliary results
Lemma A.1.
Let $X$ be a Banach space, $S\subset X$ a subspace, and $T\in\mathcal{B}(X)$ such that $T(S)$ is closed. Then $T(\overline{S})=T(S)$.
Proof.
By continuity of $T$, $T(\overline{S})\subset\overline{T(S)}$. So $T(S)\subset T(\overline{S})\subset\overline{T(S)}=T(S)$. ∎
@@
The proof is a slight modification of the argument of [12, Lemma A.2]. By separability of $\mathcal{H}$ we may a countable and dense sequence $(h_{i})_{i}$ of $\mathcal{H}$. By density of $\mathcal{D}$ we can construct a countable sequence $(d_{i}^{\prime})_{i}\subset\mathcal{D}$ dense in $\mathcal{H}$ by taking an element of $\mathcal{D}$ from the ball $B(h_{i},1/j)$, for all $i$ and $j\in\mathbb{N}$. Now, we apply Gram-Schmidt to the countable sequence $(e_{1},\ldots,e_{m},d_{1}^{\prime},d_{2}^{\prime},\ldots)\subset\mathcal{D}$ to obtain a countable orthonormal sequence $(d_{i})_{i}\subset\mathcal{D}$. Since $(e_{i})_{i=1}^{m}$ was already orthonormal, $d_{i}=e_{i}$ for $i\leq m$. Furthermore, $d_{i}^{\prime}\in\operatorname{span}{\left(d_{j},j\leq m+i\right)}$. It follows that $(d_{i}^{\prime})_{i}\subset\operatorname{span}{\left(d_{i}\right)}$ and since $(d_{i}^{\prime})_{i}$ is dense, so is $(d_{i})_{i}$. ∎
Lemma A.3.
[7, Proposition II.2.7] Let $\mathcal{H}$ be a Hilbert space. If $A\in\mathcal{B}(\mathcal{H})$, then $\lVert A\rVert=\lVert A^{*}\rVert=\lVert AA^{*}\rVert^{1/2}$.
Lemma A.4.
[18, Theorem 4.2.6] Let $\mathcal{H}$ be a Hilbert space and $A\in\mathcal{B}(\mathcal{H})$ be compact and self-adjoint. Then $\lVert A\rVert=\max\{\lvert\lambda\rvert:\ \lambda\in\sigma_{p}(A)\}$ where $\sigma_{p}(A)\coloneqq\{\lambda:\ Av=\lambda v\text{ for some }v\not=0\}$ is the point spectrum of $A$.
@@
[18, Theorem 4.3.1] Let $\mathcal{H},\mathcal{K}$ be Hilbert spaces, and $A\in\mathcal{B}(\mathcal{H},\mathcal{K})$ be compact and normal. Then $A$ is diagonalisable, that is, there exists an ONB $(h_{i})_{i}$ of $\mathcal{H}$ and an orthonormal sequence $(k_{i})_{i}$ of $\mathcal{K}$ and a sequence $(\lambda_{i})_{i}$ in $\mathbb{C}$ such that $A=\sum_{i}\lambda_{i}k_{i}\otimes h_{i}$.
@@
[7, Proposition VI.1.8] Let $\mathcal{H}$, $\mathcal{K}$ be Hilbert spaces and $A\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then $\ker{A}=\operatorname{ran}{A^{*}}^{\perp}$ and $\ker{A}^{\perp}=\overline{\operatorname{ran}{A}}$.
@@
Let $\mathcal{H}$, $\mathcal{K}$ be Hilbert spaces and $A\in\mathcal{B}_{00}(\mathcal{H},\mathcal{K})$. Then $\operatorname{rank}\left({AA^{*}}\right)=\operatorname{rank}\left({A}\right)$.
@@
Since $\operatorname{ran}{A^{*}}$ is closed, by Lemma A.7, $\operatorname{ran}{AA^{*}}=\operatorname{ran}{AP_{\operatorname{ran}{A^{*}}}}=\operatorname{ran}{AP_{\overline{\operatorname{ran}{A^{*}}}}}=\operatorname{ran}{AP_{\ker{A}^{\perp}}}=A$, where $P_{V}$ denotes the projection onto closed $V\subset\mathcal{H}$. ∎
Definition A.9.
Let $\mathcal{H}$ be Hilbert. A linear operator $A:\operatorname{dom}{A}\subset\mathcal{H}\rightarrow\mathcal{H}$ is said to be closed if its graph $\{(h,Ah):\ h\in\operatorname{dom}{A}\}$ is closed in $\mathcal{H}\oplus\mathcal{H}$.
@@
Let $\mathcal{H}$ be Hilbert, $A:\operatorname{dom}{A}\subset\mathcal{H}\rightarrow\mathcal{H}$ be closed and $B\in\mathcal{B}(\mathcal{H})$, then $AB$ is closed, $A+B$ is closed, and if $A$ is bijective, then $A^{-1}$ is closed.

### v3

@@ 5 Conclusion
Linear Gaussian inverse problems on possibly infinite-dimensional Hilbert spaces are an important kind of nonparametric inverse problem. For example, they can be used to approximate nonlinear nonparametric problems using the Laplace approximation. They often serve as the native infinite-dimensional formulation of linear inverse problems before the parameter space $\mathcal{H}$ is discretised and they are in this sense ‘discretisation independent’.
Optimal low-rank approximation of the posterior covariance for a class of losses that includes the KL divergence and the Hellinger metric, and optimal low-rank approximation of the posterior mean for the Bayes risk were studied in [31]. The analysis showed that certain matrix pencils, namely the ones defined by the Hessian and prior covariance and the prior and posterior covariance, form the central objects of study. So far, these results applied to finite-dimensional parameter spaces only.
In this work we have formulated the low-rank posterior covariance approximation problem on possibly infinite-dimensional separable Hilbert spaces. We solved this problem and derived the optimal low-rank approximations to the posterior covariance in Theorem 4.21. Equivalent conditions for its uniqueness are also given. This builds upon the finite-dimensional conclusions of [31, Section 2 and 3] for posterior covariance approximation. The resulting posterior approximation, obtained by replacing the covariance with the optimal low-rank approximation and by keeping the mean fixed, is equivalent to the exact posterior distribution, and we have shown exactly which low-rank updates of the prior covariance and precision satisfy this equivalence property in Lemma 4.5 and Proposition 4.6. Furthermore, the posterior covariance approximations are optimal for a class of losses which includes the forward and reverse KL divergences, the Hellinger metric, the Amari $\alpha$-divergences for $\alpha\in(0,1)$ and the Rényi divergences. Finally, we have shown in Proposition 3.7 that the operator pencils which proved central in the finite-dimensional analysis, are equivalent representations of the Hilbert–Schmidt operator appearing in the Feldman–Hajek theorem which quantifies similarity of Gaussian measures. For linear Gaussian inverse problems, it is therefore this operator that is central to the approximation of the posterior covariance as a low-rank update of the prior covariance. This observation is consistent with the fact that the Hilbert–Schmidt operator in the Feldman–Hajek theorem quantifies the similarity of the Gaussian prior and exact posterior.
The low-rank approximations constructed in this work provide a basis for showing the consistency of optimal low-rank covariance approximations in discretised versions of linear inverse problems. Furthermore, these approximations may be useful for the development of computationally efficient approximations of certain linear Gaussian problems. Finally, they could be used for optimal approximation of nonlinear nonparametric inverse problems.
@@ 6 Acknowledgements
The research of the authors has been partially funded by the Deutsche Forschungsgemeinschaft (DFG) Project-ID 318763901 – SFB1294. The authors thank Youssef Marzouk (Massachusetts Institute of Technology) and Bernhard Stankewitz (University of Potsdam) for helpful discussions, and Thomas Mach (University of Potsdam) for suggestions about the manuscript.
@@ Appendix A Auxiliary results
In this section we collect some auxiliary results on Hilbert spaces and bounded operators, unbounded operators and Gaussian measures.
A.1 Hilbert spaces and bounded operators
@@
The proof is a slight modification of the argument of [12, Lemma A.2]. By separability of $\mathcal{H}$ there exists a countable and dense sequence $(h_{i})_{i}$ of $\mathcal{H}$. By density of $\mathcal{D}$ we can construct a countable sequence $(d_{i}^{\prime})_{i}\subset\mathcal{D}$ dense in $\mathcal{H}$ by taking an element of $\mathcal{D}$ from the ball $B(h_{i},1/j)$, for all $i$ and $j\in\mathbb{N}$. Now, we apply Gram–Schmidt to the countable sequence $(e_{1},\ldots,e_{m},d_{1}^{\prime},d_{2}^{\prime},\ldots)\subset\mathcal{D}$ to obtain a countable orthonormal sequence $(d_{i})_{i}\subset\mathcal{D}$. Since $(e_{i})_{i=1}^{m}$ is already orthonormal, $d_{i}=e_{i}$ for $i\leq m$. Furthermore, $d_{i}^{\prime}\in\operatorname{span}{\left(d_{j},j\leq m+i\right)}$. It follows that $(d_{i}^{\prime})_{i}\subset\operatorname{span}{\left((d_{i})_{i}\right)}$ and since $(d_{i}^{\prime})_{i}$ is dense, so is $\operatorname{span}{\left((d_{i})_{i}\right)}$. ∎
Lemma A.2 ([8, Proposition II.2.7]).
Let $\mathcal{H}$ be a Hilbert space. If $A\in\mathcal{B}(\mathcal{H})$, then $\lVert A\rVert=\lVert A^{*}\rVert=\lVert AA^{*}\rVert^{1/2}$.
Lemma A.3 ([17, Theorem 4.2.6]).
Let $\mathcal{H}$ be a Hilbert space and $A\in\mathcal{B}(\mathcal{H})$ be compact and self-adjoint. Then $\lVert A\rVert=\max\{\lvert\lambda\rvert:\ \lambda\text{ is an eigenvalue of }A\}$.
@@
Lemma A.5 ([17, Theorem 4.3.1]).
Let $\mathcal{H},\mathcal{K}$ be Hilbert spaces, and $A\in\mathcal{B}(\mathcal{H},\mathcal{K})$ be compact. Then $A$ is diagonalisable, that is, there exists an ONB $(h_{i})_{i}$ of $\mathcal{H}$ and an orthonormal sequence $(k_{i})_{i}$ of $\mathcal{K}$ and a nonnegative and nonincreasing sequence $(\sigma_{i})_{i}$ such that $A=\sum_{i}\sigma_{i}k_{i}\otimes h_{i}$.
Lemma A.6 ([8, Proposition VI.1.8]).
Let $\mathcal{H}$, $\mathcal{K}$ be Hilbert spaces and $A\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then $\ker{A}=\operatorname{ran}{A^{*}}^{\perp}$ and $\ker{A}^{\perp}=\overline{\operatorname{ran}{A^{*}}}$.
@@
Let $\mathcal{H}$ and $\mathcal{K}$ be Hilbert spaces and $A\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then $\ker{AA^{*}}=\ker{A^{*}}$.
Proof.
The inclusion $\ker{A^{*}}\subset\ker{AA^{*}}$ is immediate. If $AA^{*}k=0$ for $k\in\mathcal{K}$, then $\lVert A^{*}k\rVert^{2}=\langle AA^{*}k,k\rangle=0$. Hence $A^{*}k=0$, showing the reverse inclusion holds. ∎
@@
Let $\mathcal{H}$, $\mathcal{K}$ be Hilbert spaces and $A\in\mathcal{B}_{00}(\mathcal{H},\mathcal{K})$. Then $\operatorname{ran}{AA^{*}}=\operatorname{ran}{A}$.
Proof.
Since $\operatorname{ran}{A^{*}}$ is closed, we have by Lemma A.6 that $\operatorname{ran}{AA^{*}}=\operatorname{ran}{AP_{\operatorname{ran}{A^{*}}}}=\operatorname{ran}{AP_{\overline{\operatorname{ran}{A^{*}}}}}=\operatorname{ran}{AP_{\ker{A}^{\perp}}}=\operatorname{ran}{A}$, where $P_{V}$ denotes the projection onto a closed subspace $V$ of $\mathcal{H}$. ∎
@@
Let $\mathcal{H}$ be a Hilbert space, $(e_{i})_{i}$ an orthonormal sequence, $(\delta_{i})_{i}\in\ell^{2}(\mathbb{R})$ and $T\coloneqq I+\sum_{i}\delta_{i}e_{i}\otimes e_{i}$. The following holds.
  1. (i)
    $T$ is invertible in $\mathcal{B}(\mathcal{H})$ if and only if $\delta_{i}\not=-1$ for all $i$.
  2. (ii)
    $T\geq 0$ if and only if $\delta_{i}\geq-1$ for all $i$.
  3. (iii)
    $T>0$ if and only if $\delta_{i}>-1$ for all $i$.
In cases (i) and (iii) above, the inverse of $T$ is $I-\sum_{i}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}$.
@@
Suppose that $T$ is invertible. Then $(1+\delta_{i})e_{i}=Te_{i}\not=0$ for all $i$, hence $\delta_{i}\not=-1$ for all $i$. Conversely, suppose that $\delta_{i}\not=-1$ for all $i$ and let $k\in\mathcal{H}$. Since $(\delta_{i})_{i}\in\ell^{2}(\mathbb{R})$, $\lvert(1+\delta_{i})^{-1}\rvert\leq 2$ for all $i$ large enough. Because $(\langle k,e_{i}\rangle)_{i}\in\ell^{2}(\mathbb{R})$, this implies that $\alpha\in\ell^{2}(\mathbb{R})$ where $\alpha_{i}\coloneqq(1+\delta_{i})^{-1}\langle k,e_{i}\rangle$ for all $i$. Hence $h\coloneqq\sum_{i}\alpha_{i}e_{i}\in\mathcal{H}$ and $Th=\sum_{i}(1+\delta_{i})\langle h,e_{i}\rangle e_{i}=\sum_{i}\langle k,e_{i}\rangle e_{i}=k$. This shows that $T$ is surjective. Since $T=T^{*}$, $\ker{T}=\operatorname{ran}{T}^{\perp}=\{0\}$ by Lemma A.6, showing that $T$ is injective, which proves (i). If $T\geq 0$, then $1+\delta_{i}=\langle Te_{i},e_{i}\rangle\geq 0$, i.e. $\delta_{i}\geq-1$, for all $i$. Conversely, if $\delta_{i}\geq-1$ for all $i$, then $\langle Th,h\rangle=\sum_{i}(1+\delta_{i})\langle h,e_{i}\rangle^{2}\geq 0$. This proves (ii), and replacing “$>$” by $\geq$”, also (iii).
To compute the inverse of $T$, note that $\frac{\delta_{i}}{1+\delta_{i}}\leq 2\delta_{i}$ for all $i$ large enough, by the hypothesis that $(\delta_{i})_{i}\in\ell^{2}(\mathbb{R})$. Thus, $\frac{\delta_{i}}{1+\delta_{i}}\rightarrow 0$ and $\sum_{i}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}$ is well-defined in $\mathcal{B}(\mathcal{H})$. For $h\in\mathcal{H}$, we have by direct computation,
$=\sum_{i}\left(1-\frac{\delta_{i}}{1+\delta_{i}}+\delta_{i}-\frac{\delta_{i}^{2}}{1+\delta_{i}}\right)\langle h,e_{i}\rangle e_{i}$
$=\sum_{i}\langle h,e_{i}\rangle e_{i}=h.$
Similarly,
$\left(I-\sum_{i}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}\right)\left(I+\sum_{i}\delta_{i}e_{i}\otimes e_{i}\right)h=$ $\left(I-\sum_{i}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}\right)\sum_{i}\left(1+\delta_{i}\right)\langle h,e_{i}\rangle e_{i}$
$=$ $\sum_{i}\left(1+\delta_{i}-\frac{\delta_{i}}{1+\delta_{i}}(1+\delta_{i})\right)\langle h,e_{i}\rangle e_{i}.$
@@
Let $\mathcal{H},\mathcal{K}$ be Hilbert spaces. Suppose $A_{1},A_{2}\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then the following are equivalent:
  1. (i)
    $\operatorname{ran}{A_{1}}\subset\operatorname{ran}{A_{2}}$,
  2. (ii)
    There exists $C>0$ such that $\{A_{1}h:\ \lVert h\rVert\leq 1\}\subset\{A_{2}h:\ \lVert h\rVert\leq C\}$,
  3. (iii)
    There exists $C>0$ such that $\lVert A_{1}^{*}k\rVert\leq C\lVert A_{2}^{*}k\rVert$ for all $k\in\mathcal{K}$.
@@
See [11, Proposition B.1(i)] and its proof. ∎
Definition A.11 ([8, Definition VIII.3.10]).
Let $\mathcal{H}$ be a Hilbert space. We say that $W\in\mathcal{B}(H)$ is a ‘partial isometry’ if $W$ is an isometry on $\ker{W}^{\perp}$. We call $\ker{W}^{\perp}$ the ‘initial space’ of $W$ and $\operatorname{ran}{W}$ the ‘final space’ of $W$.
Recall from Section 1.5 that $\lvert A\rvert\coloneqq(A^{*}A)^{1/2}$ for $A\in\mathcal{B}(\mathcal{H})$. For a proof of the following, see e.g. [8, VIII.3.11].
Lemma A.12 (Polar decomposition).
Let $\mathcal{H}$ be a Hilbert space and $A\in\mathcal{B}(\mathcal{H})$. There exists a partial isometry $W\in\mathcal{B}(\mathcal{H})$ with initial space $\ker{A}^{\perp}$ and final space $\ker{A^{*}}^{\perp}$ such that $A=W\lvert A\rvert$.
Lemma A.13.
Let $\mathcal{H}$ be a Hilbert space and $A,B\in\mathcal{B}(\mathcal{H})$ be injective with $\operatorname{ran}{AA^{*}}$ dense. If $AA^{*}=BB^{*}$, then there exists a Hilbert space isomorphism $Q\in\mathcal{B}(\mathcal{H})$ such that $B=AQ$.
Proof.
We first note that $\operatorname{ran}{A}$ is dense, since $\operatorname{ran}{AA^{*}}=A(\operatorname{ran}{A^{*}})\subset A(\overline{\operatorname{ran}{A^{*}}})=A(\ker{A}^{\perp})=A(\mathcal{H})=\operatorname{ran}{A}$ by Lemma A.6 and $\ker{A}=\{0\}$. Since $AA^{*}=BB^{*}$, also $\operatorname{ran}{BB^{*}}$ and $\operatorname{ran}{B}$ are dense. Now, by the polar decomposition applied to $A^{*}$ and $B^{*}$, c.f. Lemma A.12, there exist $W_{1},W_{2}\in\mathcal{B}(\mathcal{H})$ such that $A^{*}=W_{1}\lvert A^{*}\rvert$, $B^{*}=W_{2}\lvert B^{*}\rvert$. Here, $W_{1}$ is an isometry on $\ker{A^{*}}^{\perp}$ with $\operatorname{ran}{W}_{1}=\ker{A}^{\perp}$, and similarly, $W_{2}$ is an isometry on $\ker{B^{*}}^{\perp}$ with $\operatorname{ran}{W_{2}}=\operatorname{ran}{B}^{\perp}$. Since $\ker{A}=\{0\}$ by assumption, it follows that $W_{1}$ is surjective. Since $\operatorname{ran}{A}^{\perp}=\ker{A^{*}}=\{0\}$ by assumption and Lemma A.6, it follows that $W_{1}$ is an isometry on all of $\mathcal{H}$. Hence $W_{1}$ is a surjective isometry on $\mathcal{H}$, that is a Hilbert space isomorphism. Similarly, $W_{2}$ is a Hilbert space isomorphism, and therefore so is $W_{2}W_{1}^{-1}$. Now, $AA^{*}=BB^{*}$ implies $\lvert A^{*}\rvert=\lvert B^{*}\rvert$. Thus, $B^{*}=W_{2}\lvert B^{*}\rvert=W_{2}\lvert A^{*}\rvert=W_{2}W_{1}^{-1}W_{1}\lvert A^{*}\rvert=W_{2}W_{1}^{-1}A^{*}$. We conclude that $B=AQ$, where $Q\coloneqq(W_{1}W_{2}^{-1})^{*}\in\mathcal{B}(\mathcal{H})$ is a Hilbert space isomorphism. ∎
A.2 Unbounded operators
For $A\in\mathcal{B}(\mathcal{H})$, we denote by $A^{\dagger}$ the Moore–Penrose inverse of $A$, also known as the generalised inverse and pseudo-inverse of $A$, c.f. [13, Definition 2.2], [11, Section B.2] or [17, Definition 3.5.7]. It holds that $A^{\dagger}$ is bounded if and only if $\operatorname{ran}{A}$ is closed, c.f. [13, Proposition 2.4]. If $A$ is injective, then $A^{\dagger}=A^{-1}$ on $\operatorname{ran}{A}$.
Lemma A.14.
Let $\mathcal{H}$ and $\mathcal{K}$ be Hilbert spaces. Suppose $A_{1},A_{2}\in\mathcal{B}(\mathcal{H},\mathcal{K})$. If $\operatorname{ran}{A_{1}}\subset\operatorname{ran}{A_{2}}$, then there exists $C>0$ such that $\lVert A_{2}^{\dagger}k\rVert\leq C\lVert A_{1}^{\dagger}k\rVert$ for all $k\in\operatorname{ran}{A_{1}}$.
The proof is a modification of the arguments in the proof of [11, Proposition B.1(ii)].
Proof.
Let us first assume that $A_{2}$ is injective, so that $A_{2}^{\dagger}=A_{2}^{-1}$ on $\operatorname{ran}{A_{2}}$. We must show that there exists $C>0$ such that $\lVert A_{2}^{-1}k\rVert\leq C\lVert A_{1}^{\dagger}k\rVert$ for all $k\in\operatorname{ran}{A_{1}}$. We shall obtain a contradiction by supposing no $C>0$ exists such that $\lVert A_{2}^{-1}k\rVert\leq C\lVert A_{1}^{\dagger}k\rVert$ for all $k\in\operatorname{ran}{A_{1}}$. That is, we suppose that for each $m\in\mathbb{N}$ there exists $k^{m}\in\operatorname{ran}{A_{1}}$ such that $\lVert A_{2}^{-1}k^{m}\rVert>m\lVert A_{1}^{\dagger}k^{m}\rVert$. Since $k^{m}\in\operatorname{ran}{A_{1}}$ and $\operatorname{ran}{A_{1}}\subset\operatorname{ran}{A_{2}}$, there exist $\tilde{h}_{1}^{m},\tilde{h}_{2}^{m}\in\mathcal{H}$ such that $\tilde{h}_{1}^{m}=A_{1}^{\dagger}k^{m}$ and $\tilde{h}_{2}^{m}=A_{2}^{-1}k^{m}$. Thus, $A_{1}\tilde{h}^{m}_{1}=A_{2}\tilde{h}^{m}_{2}=k^{m}$. Define $h_{i}^{m}\coloneqq\tilde{h}_{i}^{m}/\lVert\tilde{h}_{1}^{m}\rVert$, $i=1,2$. Then $\lVert h_{1}^{m}\rVert=1$ for all $m$ and $\lVert h_{2}^{m}\rVert\rightarrow\infty$ as $m\rightarrow\infty$. On the one hand, for every $C>0$ there exists $M\in\mathbb{N}$ such that $A_{2}h_{2}^{m}\not\in\{A_{2}h:\ \lVert h\rVert\leq C\}$ for all $m>M$, by injectivity of $A_{2}$. On the other hand, $A_{1}h^{m}_{1}=k^{m}/\lVert\tilde{h}_{1}^{m}\rVert=A_{2}h^{m}_{2}$, hence $A_{2}h^{m}_{2}\in\{A_{1}h:\ \lVert h\rVert\leq 1\}$ for all $m$. By Lemma A.10, $A_{2}h^{m}_{2}\in\{A_{2}h:\ \lVert h\rVert\leq C\}$ for all $m$ and for some $m$-independent constant $C>0$, which is a contradiction.
Now let $A_{2}\in\mathcal{B}(\mathcal{H},\mathcal{K})$ be arbitrary. The subspace $\ker{A}_{2}^{\perp}\subset\mathcal{H}$ is closed and therefore a Hilbert space with respect to its subspace topology. Let us denote the restriction of $A_{2}$ to $\ker{A_{2}}^{\perp}$ by $\tilde{A}_{2}\in\mathcal{B}(\ker{A_{2}}^{\perp},\mathcal{K})$. Then $\tilde{A}_{2}$ is injective and satisfies $\operatorname{ran}{\tilde{A}_{2}}=\operatorname{ran}{A_{2}}$. By construction of the Moore–Penrose inverse, $A_{2}^{\dagger}k=\tilde{A}_{2}^{-1}k\in\mathcal{H}$ for $k\in\operatorname{ran}{\tilde{A}}_{2}=\operatorname{ran}{A_{2}}$. By the hypothesis $\operatorname{ran}{A}_{1}\subset\operatorname{ran}{A}_{2}$, we have $A_{2}^{\dagger}k=\tilde{A}_{2}^{-1}k\in\mathcal{H}$ for $k\in\operatorname{ran}{A_{1}}$. From the previous part of the proof we can then conclude the existence of $C>0$ such that $\lVert A_{2}^{\dagger}k\rVert=\lVert\tilde{A}_{2}^{-1}k\rVert\leq C\lVert A_{1}^{\dagger}k\rVert$ for all $k\in\operatorname{ran}{A_{1}}$. ∎
Definition A.15 ([8, Definition X.1.3]).
Let $\mathcal{H}$ be a Hilbert space. A linear operator $A:\operatorname{dom}{A}\subset\mathcal{H}\rightarrow\mathcal{H}$ is said to be closed if its graph $\{(h,Ah):\ h\in\operatorname{dom}{A}\}$ is closed in $\mathcal{H}\oplus\mathcal{H}$.
Lemma A.16.
Let $\mathcal{H}$ be a Hilbert space, $A:\operatorname{dom}{A}\subset\mathcal{H}\rightarrow\mathcal{H}$ be closed and $B\in\mathcal{B}(\mathcal{H})$. Then,
  1. (i)
    $AB$ is closed,
  2. (ii)
    $A+B$ is closed,
  3. (iii)
    if $A$ is also injective, then $A^{-1}:\operatorname{ran}{A}\subset\mathcal{H}\rightarrow\operatorname{dom}{A}\subset\mathcal{H}$ is closed.
Proof.